### 16 Feb 2024 v3

# Calibration and Correctness of Language Models for Code

Abstract.
Machine learning models are widely used but can also often be wrong. Users would benefit from a reliable indication of whether a given output from a given model should be trusted, so a rational decision can be made whether to use the output or not. For example, outputs can be associated with a confidence measure; if this confidence measure is strongly associated with likelihood of correctness, then the model is said to be well-calibrated. In this case, for example, high-confidence outputs could be safely accepted, and low-confidence outputs rejected.
Calibration has so far been studied in mostly non-generative (e.g., classification) settings, especially in Software Engineering. However, generated code can quite often be wrong: Developers need to know when they should e.g., directly use, use after careful review, or discard model-generated code; thus Calibration is vital in generative settings. However, the notion of correctness of generated code is non-trivial, and thus so is Calibration. In this paper we make several contributions. We develop a framework for evaluating the Calibration of code-generating models. We consider several tasks, correctness criteria, datasets, and approaches, and find that by and large generative code models are not well-calibrated out of the box. We then show how Calibration can be improved, using standard methods such as Platt scaling. Our contributions will lead to better-calibrated decision-making in the current use of code generated by language models, and offers a framework for future research to further improve calibration methods for generative models in Software Engineering.
LLMs, Calibration, Confidence Measure
copyright: none
1. Introduction
Generative large language models (LLMs) are now widely-used for code completion in IDEs; in fact, LLMs have now reportedly become integrated into a wider range of software development activities including code review, code summarization, program repair, etc.. For example, Google’s Didact system (researchLargeSequence, 35), supports debugging & repair, code review, and editing. However, LLMs come with documented risks: they might generate known buggy code (10174227, 25), or code with known & risky vulnerabilities (asare2022github, 3, 52). The potential for bad code being generated by LLM “copilots” (lo2023trustworthy, 40) is a widespread and growing risk to the software infrastructure that most modern societies rely on.
Nevertheless, current trends suggest that these LLM-based tools are becoming ever-more integrated into software development practices, potentially ushering in a future where most code is machine-written. The motivation for our work is twofold: first, given that LLM-generated outputs could be wrong, humans who use such outputs would benefit from a clear and reliable indication of confidence that the generated code is correct. Second, such indications of confidence, when reliable and well-calibrated, can support sound engineering responses to possible quality issues, and thus improve the safety and reliability of the AI4SE ecosystem. Thus, for example, if an LLM offers a code suggestion, such a suggestion should ideally be accompanied by a numerical “confidence” signal, (e.g., a probability) indicating how likely the generated code is to match intent correctly. This signal should be well-calibrated: code suggestions with a high confidence score should usually be very good, and we should be able to reason about the likelihood and risks in a principled way. While calibration has been used in other settings (e.g. classically in weather prediction), our focus here is to empirically study the calibration of generative large language models across practical software engineering settings, such as line-level code completion, function synthesis, and code-repair. When available, a well-calibrated measure would facilitate better risk-management in the development process, and eventually lead to higher code quality.
For example, a development team might reasonably adopt a policy that: a) generated code suggestions associated with a high confidence in lower risk domains could be reviewed lightly and quickly accepted; b) suggestions with a medium confidence value should be reviewed more carefully before acceptance; and c) those with a low confidence value should be simply rejected or the prompt should be adjusted. In this paper, we make the following contributions:
  • We explore the notion of calibration for confidence measures of LLM-generated code.
  • We cover multiple generative tasks, datasets, and confidence measures, providing wide perspective on the calibration of code generation.
  • We evaluate calibration using two different measures of correctness, including exact-match, and correctness-modulo-testing.
  • We evaluate calibration of both naturally-occurring code, from active GitHub projects, and code drawn from artificial, coding-challenge datasets.
We conclude our paper with some thoughts on future directions.
2. Background
We begin first with some basic concepts, and then present our research agenda.
2.1. Basic Concepts
We introduce the idea of calibration, and then discuss its potential value when using large language models for coding.
2.1.1. Calibration:
Calibration has deep roots in prediction problems like weather forecasting. Suppose a weather model predicts a 70% confidence (probability) of rain the next day. Is this confidence score helpful? As discussed in (dawid1982well, 15), for a well-calibrated model, out of 100 days where the model predicted rain with 70% confidence, then in roughly 70 of those days it should have actually rained. That is, a model is well-calibrated if the model’s confidence in a given output, is quite close to the empirical relative frequency (likelihood) with which the output is actually correct. Good calibration supports further decision-making based on the confidence of a model. Poorly-calibrated, overconfident model—that often produces wrong output with high confidence—may lead to poor decisions; as might a diffident one that often outputs low-confidence when it is correct. A well-calibrated model informs the user exactly what the expectation of the prediction is. Thus, at varying degrees of confidence, a well-calibrated model gives a user the opportunity to respond in a calibrated way: at 20% confidence of rain, one might take a hat; at higher chances, one might carry an umbrella; at even higher chances, one might choose to take a bus rather than walk, etc. The idea of calibration has recently been a topic of interest in natural language processing. From an earlier work by (jiang2021can, 27): given a model $M$, an input $X$ and true, expected correct output $Y$, a model output $M(X)=\hat{Y}$, (note that we won’t always have $Y=\hat{Y}$) and a output probability $P_{M}(\hat{Y}\,|\,X)$ provided by the model, a perfectly calibrated model satisfies the following condition (Adapted from (jiang2021can, 27)):
[…]
In other words, if we have perfectly-calibrated confidence ${p}$ (the model’s calculated probability of its prediction that the output is $\hat{Y}$), then this value is exactly the same as the empirical fraction p of the cases where the actual output $Y$ is correctly matches the prediction $\hat{Y}$. These notions of confidence, and correctness are important in this paper, and the reader may find the above equation a useful point of reference. The ${p}$ refers to the model confidence; the p refers to the correctness (more precisely, the fraction of predictions that are correct and match intent).
Usually, the level of calibration is imperfect; the yellow p above is typically not exactly the same as the first pink ${p}$. One empirically estimates the value of the yellow p from experimental samples, and measures the difference to the actual provided value. There are various measures of this difference, including Brier Score (brier1950verification, 9) and $ECE$ (Expected Calibration Error) (Pakdaman_Naeini_Cooper_Hauskrecht_2015, 48), which we describe below.
2.1.2. Why Calibration matters for Code
LLMs do make mistakes (10174227, 25), despite their wide use and popularity (vaithilingam2022expectation, 59, 71). To manage the risks of accepting bad code from an LLM, developers could use (lo2023trustworthy, 40, 19) a well-calibrated signal of confidence ${p}$ associated with generated code, for example. A well-calibrated high-value of ${p}$ would suggest a high empirical probabilty p that the code is correct. This sort of reliable signal of confidence in generated-code correctness supports better decision-making and proper use of code review resources, with well-placed trust and ready acceptance for code produced by the model with high-confidence of correctness, and appropriately careful review of code produced with less confidence. A poorly calibrated model may lead to either unnecessary review of likely correct code, or ill-advised acceptance of likely incorrect code. It should be noted here that good calibration allows more nuanced, effective quality-control (Q-C) decisions, beyond simple binary decisions (e.g. carefully review each token of generated code vs. just use it). For example, with a well-calibrated model, over 90% confidence might indicate that the generated code can be accepted with minimal review; below 10% confidence might indicate poor-quality code that should just be rejected; at 30% confidence the code might be subjected careful review, perhaps by more than one developer; at 70% perhaps the user themselves could review it, and accept it, and so on. This decision making also ties into reasoning about the risks of the domain. Such a well-calibrated quality-control process has been used in medicine e.g., for elder-care (nierman2001outcome, 44), and for decision-making in cancer-care (schwarz2019guess, 53). While all these decision points should also consider the relative cost of these different Q-C decisions, a well-calibrated confidence signal is an important factor. Overall, given the cost & consequences of properly addressing software quality, and the potential benefits of LLM-generated code, a well-calibrated confidence signal is highly desirable.
2.2. Research Agenda
In software engineering, LLMs are perhaps mostly widely used for code suggestion/completion. Other tasks include code synthesis, and program repair. We therefore wish to study calibration for tasks involving code-generation. While there are many other applications of LLMs, this paper only considers code synthesis, completion, and repair.
RQ 1. How well are language models calibrated for three tasks: Function Synthesis, Line-level Code Completion, and Program Repair?
Calibration requires a “confidence measure” (see § 3.1.1). We evaluate calibration for several confidence measures. First, we assess probabilities directly produced by the model for a given generated output (intrinsic calibration). Alternatively, we can instruct an LLM to “reflectively” consider its own outputs, and produce an assessment (reflective calibration).
RQ 2. How well do intrinsic calibration approaches work?
RQ 3. How well do reflective calibration approaches work?
Third, there are various approaches to correctness, with different characteristics. We evaluate calibration of language models of code with respect to two approaches ( § 3.1.1).
RQ 4. How calibrated are LLMs for code with respect to “correctness-modulo-testing” and “exact-match” notions of correctness?
Finally, we consider the value of rescaling confidence scores produced by models.
RQ 5. Does rescaling improve calibration performance?
3. Research Methodology
As discussed earlier, our primary focus is on generative model calibration rather than accuracy. We consider three generative tasks i.e., function synthesis, line-level code completion, and program repair, where generative LLMs are directly applicable and widely-used (e.g., completion in Copilot). In this section, we will discuss the tasks, datasets, models, and methodology of our approach.
3.1. Tools & Measures
We now describe our experimental tools & measures.
3.1.1. Code Correctness
When evaluating calibration, we need a notion of correctness. For (non-generative) models outputting labels, classes, True/False, etc. correctness is simply an exact-match with ground-truth correct label. Exact-match could also be viewed as a notion of correctness for code, e.g., with defect repair, where there is a known, incorrect “buggy” version, and a known “fixed” version. If generated code matches the fixed version given an appropriate prompt, it is “correct”; otherwise, it is not. However, this approach is overly strict; the generated code might not be an exact-match, and yet still be “correct” based on test-passing: generated code not exactly matching some given “correct” version, might still pass the given tests. Other notions of correctness exist: code-review, formal verification, etc. We use test cases provided with the code as our preferred indication of correctness; tests are widely used, and are easily automatable.
Unlike exact-match correctness test-passing correctness offers the advantage of admitting different semantically identical forms. However, test-passing correctness may be too permissive, or even wrong: test cases maybe insufficient or incorrect [^1] Recently, there has been a paper (liu2023code, 39) accepted to NeurIPS, reporting inadequacies in the test sets of HumanEval; in addition tests can be “flaky” (luo2014empirical, 42): the same test, on the same code, can sometimes pass and sometimes fail. In conclusion, both test-passing and exact-match afford automation, and are thus both appealing in an experimental setting such as ours; but both have limitations (see discussion of Table 4 below).
In this work, we studied both: we report test-passing correctness wherever tests are available, and when they are not, we report exact-match correctness. For the cases where tests were available, the exact-match results are presented in the appendix (see Table 7).
3.1.2. Confidence measures
We usually calculate a confidence measure (or probability) $p$, associated with a generated output code text $C$. We consider two categories of measures: intrinsic probability, which is calculated by the generating language model itself, and reflective probability, obtained by re-invoking the model, prompting it to estimate how confident it is in the correctness of of the code it had just generated (See Figure 1). Our measures include:
Figure 1. Prompts for Verbalized Self-Ask and Question Answering logit.
Average Token Probability:
(Intrinsic, $p_{avg}$) Given a candidate output sequence $\mathrm{T}$ of tokens
$\tau_{1},\tau_{2},\ldots\tau_{n}$, we collect the associated (auto-regressive) model probabilities [^2] The models can provide a logit for a selected token, given previous tokens; the logit can be exponentiated to a probability. $p(\tau_{i})$ of each token $\tau_{i}$, and then average these probabilities (arithmetic mean) to yield $p_{\text{avg}}(\mathrm{T})$
[…]
(2) $p_{\text{avg}}({\mathrm{T}})=\frac{1}{n}\displaystyle\sum_{i=1}^{n}p(\tau_{i})$
Generated Sequence Probability:
(Intrinsic, $p_{tot}$) The probability of the full generated sequence is calculated as the product of all the probabilities:
[…]
(3) $p_{tot}({\mathrm{T}})=\displaystyle\prod_{i=1}^{n}p(\tau_{i})$
Verbalized Self-Ask:
(lin2022teaching, 38, 70, 58) (Reflective, $p_{v}$) We prompt the model to reflect jointly on the prompt, and the model-generated code, and then output a numerical value of its confidence in the generated code. An example is shown in Figure 1, on the left hand side. We use string-matching heuristics to handle percentages appearing different forms (e.g. “80%”, “80.00%”, “The probability is about 80.”, etc). For the line completion task, we experimented with a task-specific variation of the prompt, but found no improvement in rescaled calibration scores. Some extra logic is implemented for when the model fails to output a probability  [^3] Discussed further in appendix Understanding effects of verbalized retry failures.
Question Answering Logit:
(kadavath2022language, 30) (Reflective, $p_{B}$ and $p_{NB}$) In this case, rather than prompting for a numerical score (as above), we prompt for a TRUE or FALSE answer. The probability associated with the TRUE token is taken to be the confidence measure. [^4] We transform raw Logits from the output layer of the LLM into probabilities using the softmax function. Additionally we extend this approach using normalization: the model can assign probability mass to multiple possible expressions of TRUE or FALSE or even other (e.g, “ True”, “ true”, “ ”, ). Thus when extending to the normalized form (Ask T/F N), we take the fraction of probability mass “True” assigned between only “True” and “False”. As GPT-3.5-Turbo-Instruct does not provide log-probabilities of the input tokens, we took the five most likely tokens for the first two completed tokens and selected the token with the maximum probability containing “true” or “false” (case insensitive), respectively.
Other Methods:
(Mixed) Other methods have been proposed (zhang2023coder, 68, 32). Self-consistency could also be used (wang2022self, 60, 1) could be another option. These are left for future research.
Our experiments consider the first four above: average token probability, generated sequence probability, verbalized self-evaluation, and question answering logit. In addition, as a baseline, we also used the length of the generated sequence. The length baseline is calculated based on the number of characters in the generated sequence, scaled such that 00 is the shortest value in the dataset, and $1$ is the longest.
For reflective measures, we expected that a code generation model should perform well (and provide well-calibrated confidence scores for correctness): first, for synthesis, given just a good natural language description (without tests), of the desired function; second, for completion & bug-fixing, given the surrounding context. In both settings, we measure correctness using available hidden test cases. We also note that including a failing test in an LLM prompt is not common experimental practice for the tasks we analyze. (chen2021evaluating, 12, 4, 26).
3.1.3. Measures of Calibration
Whenever there are measures of a model’s confidence in its generated output, and a reliable way of determining if the output is correct, one can use standard measures of calibration. Below we will discuss in more detail: confidence & correctness measures, issues in measuring them, and standard calibration measures. Calibration measures are conceptually based on the empirical calibration plot (Also, referred to as Reliability Plot.), which plots confidence vs correctness. For example, 1(a) is a plot based on bucketing experimental samples of a given confidence measure.
(a) Well Calibrated
(b) Poorly Calibrated
Figure 2. Sample calibration plots demonstrating well- vs. poorly- calibrated.
To obtain larger samples of outputs, and thus more robust empirical estimates of the correctness probability p (from Equation 1), one typically buckets outputs with similar model-estimated confidences (probabilities $\hat{p}$quation 1).
Two empirical calibration plots illustrating this method are shown in Figure 21(a) is for token-level code completions from Codex; it shows the observed proportion of exact-matches (y-axis) vs the predicted probability i.e., confidence as per the language model, based on bucketing observations into subsets $S_{1},S_{2},\ldots S_{n}$. Here, there are $n=10$ buckets, equally spaced by confidence measure; for each bucket, the height of the bar indicates the proportion (value $\in[0,1]$) of correct samples in the bucket. The closer the bars in each $S_{i}$ are to the diagonal line, the better the calibration.
We note here that Codex is a large model, well-trained on the task of token-level completion; as expected, it is both well-calibrated and generally correct on the simple token-level completion task. However, for notions of correctness farther from the training objective, such as line-level code completion and test-passing, Codex’s [intrinsic] probability may not be as well-calibrated. An example where the confidence is not well-calibrated is GPT-3.5 verbalized confidence on DyPyBench (No Docstring) with regard to “all-passed” notion of correctness. This is presented in 1(b).
We use two measures of (mis)calibration: Brier Score (brier1950verification, 9) and $ECE$ (Pakdaman_Naeini_Cooper_Hauskrecht_2015, 48). Both measure the deviation from perfect calibration. As before (following (guo2017calibration, 20)), we assume a model $M$, input $X$, actual desired output $Y$, and model prediction $M(X)=\hat{Y}$.
In our case, both $Y$ and $\hat{Y}$ are code, rather than a single label. We later discuss notions of equivalence. For convenience, we denote the model’s confidence as $\hat{P}$, which we want to be well-calibrated. Thus the “perfect calibration equation” requires that:
[…]
(4) $P(\hat{Y}=Y\mid\hat{P}=p)=p,~{}~{}~{}~{}\forall p\in[0,1]$
[…]
(5) $\text{corr}(S_{i})=\frac{1}{\mid S_{i}\mid}\displaystyle\sum_{x_{i}\in S_{i}}\mathds{1}(\checkmark M(x_{i}))\qquad\mathrm{and}\qquad\text{conf}(S_{i})=\frac{1}{\mid S_{i}\mid}\displaystyle\sum_{x_{i}\in S_{i}}\hat{p}_{i}$
where the model generates the code $M(x_{i})$ with confidence (probability) $\hat{p}_{i}$, and $\checkmark M(x_{i})$ indicates that the generated code is correct, as per the operative experimental definition. For a perfectly calibrated $M$, we would have $\text{corr}(S_{i})=\text{conf}(S_{i})\quad\forall i\in 1\ldots m$. In practice, we observe deviations from this ideal. Expected Calibration Error ($ECE$(Pakdaman_Naeini_Cooper_Hauskrecht_2015, 48) is a typical measure of calibration, calculated as the weighted average of these deviations.
[…]
(6) $ECE=\displaystyle\sum_{i=1}^{m}\frac{\mid S_{i}\mid}{\mid T\mid}\lvert\text{corr}(S_{i})-\text{conf}(S_{i})\rvert$
$ECE$ is intuitive, but can mislead; as seen below, a naive predictor whose confidence is always the base rate would yield a deceptively low $ECE$ value. An alternative measure of calibration, which avoids this issue is the Brier Score ${\mathcal{B}}$ (brier1950verification, 9); it
[…]
(7) $\mathcal{B}=\frac{1}{\mid T\mid}\displaystyle\sum_{i=1}^{\mid T\mid}(\hat{p}_{i}-\mathds{1}(\checkmark M(x_{i})))^{2}$
One achieves an optimal Brier of $\mathcal{B}\approx 0$ when confidently estimating $\hat{p}_{i}\approx 1$ when the code is correct, and estimating $\hat{p}_{i}\approx 0$ when the code is incorrect for each sample.
[…]
(8) $\mathcal{B}_{ref}=p_{r}(1-p_{r})$
[…]
(9) $SS=\frac{\mathcal{B}_{ref}-\mathcal{B}_{actual}}{\mathcal{B}_{ref}}$
Positive $SS$ (perfect score = 1.0) indicates improvement over baseline $\mathcal{B}_{ref}$; negative indicates making predictions worse than the baseline. Small positive values of $SS$ can sometimes indicate good skill; for example, the Deutsche Wetterdienst (German weather forecasting service) considers 0.05 Skill Score to be a minimum threshold for a good forecast quality  (dwdWetterKlima, 62). As another data point, the American data journalism site 538 reports around a 0.13 skill in forecasting World Cup games (fivethirtyeightGoodFiveThirtyEight, 24), which is in the range of what we observe in our experiments for best case code generation by LLMs; but these LLM performance numbers are just a starting point, and can be expected to improve in the future. $ECE$ (Equation 6) and Brier Score (Equation 7) serve slightly different purposes: the Brier Score is calculated for each sample, and measures both the ability to correctly discriminate output categories, and calibration of the output probability. The $ECE$ measures just calibration; but it can be misleadingly low, as noted above for the unskilled predictor. Additionally binning must be done carefully, since it can affect $ECE$ scores (nixon2019measuring, 47).
3.1.4. Rescaling Approaches
Machine learning models are not always calibrated. Guo et al (guo2017calibration, 20) discuss ways of rescaling probability estimates to better match observations. A common approach is Platt scaling (Platt1999ProbabilisticOF, 50), where a logistic regression is fit to the logit values of the prediction i.e., the $\ln$ of the measured confidence probability. This optimizes two parameters, a linear scaling multiplier and a bias i.e., intercept, that shifts the value.
To reduce the likelihood that the scaling overfits & skews our results, we rescale over five folds; i.e., we fit a logistic regression on a random $\nicefrac{{4}}{{5}}$ of data and apply it to $\nicefrac{{1}}{{5}}$ of data, before sliding over and doing each combination of $\nicefrac{{4}}{{5}}$.
Besides Platt scaling, temperature rescaling has also been used (desai2020calibration, 16, 29, 20): this approach applies a scalar multiplier on the logits representing each class e.g., a multiclass image classifier. In our binary confidence case, this has similar expressivity to Platt scaling without an intercept. Other approaches include histogram binning (Zadrozny2001ObtainingCP, 63), isotonic regression (Zadrozny2002TransformingCS, 64), inter alia (guo2017calibration, 20, 34). These approaches are more parameterized; given the data limitations in our experimental setup (e.g., a few hundred examples in function synthesis) pose higher risk of overfitting.
As we discuss in section 4, Platt scaling does improve calibration, with some caveats.
3.1.5. Receiver Operating Characteristic vs. Calibration
Metrics such as Brier Score and $ECE$ are designed for calibration. However, one might be interested in treating correctness of generated code as a typical classification task, and thus apply familiar metrics like accuracy or Area under Curve of the Receiver Operating Characteristic (AUC-ROC abbr. AUC). AUC is typically a favored metric as it is parameter free, meaning that one does not choose a threshold for when to label values as incorrect or correct from the score. AUC and Scaled Skill Score (discussed further below) serve a similar purpose, and are indeed highly correlated (appendix Figure 7). However, there are differences between AUC and calibration. Notably, well calibration measures aid in more granular decision processes, and that it is possible for measures with similar AUCs to have different calibrations and thus different $ECE$. This aspect is discussed further in section 5. We note how measures with roughly equivalent AUC can have noticeably different $ECE$ values (see appendix Figure 8). Thus AUC does not tell the whole story.
Task Dataset Dataset Size
Correctness
Measure
Confidence Measure
Calibration
Metric
Function synthesis HumanEval 164 Test-passing Correctness Average Token Probability, Generated Sequence Probability, Verbalized Self-Evaluation, Question Answering Logit Brier Score, $ECE$
MBBP 974
Line-level Completion DyPyBench 1,988 Test-passing Correctness, EM
Program Repair Defects4J 120
ManySStubs4j 3,000 Exact-Match (EM)
Table 1. List of tasks with associated datasets and evaluation measures.
Figure 3. Prompt and model output for the tasks while calculating confidence measure based on Average Token Probability and Generated Sequence Probability.
3.2. Tasks & Dataset
3.2.1. Function Synthesis
This task aims to synthesize/generate Python functions from “Docstrings”. [^5] Docstrings are code comments that explain the code’s purpose and usage. Further discussed in § 3.2.2. Correctness is determined by functional testing.
Figure 3a shows a sample prompt and model output. We use HumanEval (chen2021evaluating, 12) and MBPP (austin2021program, 4) datasets. One caveat: the samples in these datasets largely constitute artificial problems, specifically assembled to test the code-synthesis capacity of large language models; performance results (both accuracy and calibration) over these datasets may not necessarily generalize to real-world software development. Even so, these datasets provide a valuable datapoint for assessing model calibration.
The HumanEval dataset is per se structured as function synthesis. However, MBPP is natively a file generation problem: for each sample, we are given a prompt, the reference solutions might include imports (approximately 20% of problems) or small helper classes or functions (approximately 5% of problems). We restructure all problems into a function synthesis task by placing the prompt inside the tested method as a Docstring (but possibly allowing the model to see other context). This was done for two reasons. First, it makes it comparable to HumanEval and the task of function synthesis. Second, it makes it more feasible for zero-shot auto-regressive models, as we can truncate generation as soon as the function ends. Without truncation, a normal LLM tends to continue generating arbitrary code until the maximum length has been reached, which can introduce bugs unrelated to solving the problem. Additionally we exclude approximately 75 problems where the reference solution fails to pass the provided test cases [^6] Due to either buggy reference code/tests, or possibly missing environment/networking/compute-time requirements.
3.2.2. Line-level Code Completion
Code completion is a well-studied task in Software Engineering, and widely deployed in IDEs, with tools like GitHub Copilot (chen2021evaluating, 12). This task can be investigated at both the token and line levels (izadi2022codefill, 23, 33, 41). Performance on this task provides a good indication of model’s ability to generate a sequence of code tokens, conditioned on a particular context (viz., “auto-regressive” generation); usually the preceding sequence of prompt tokens. In real development settings, developers may not always require the complete function from the model. Instead, generating the next several tokens is often helpful enough. Full-function code synthesis is more challenging than small-scale completion; moreover, partial code completion is already in wide use, and has been evaluated using real software development data (ziegler2022productivity, 71, 57); however, calibration for code-completion, which is the most widely-deployed task for LLMs, has so far not been evaluated in detail. The current decoder-only GPT models (brown2020language, 10, 12) are already trained to generate the next token at low average cross-entropy given all the prior tokens, following the condition $p(token\ |\ prior\ tokens)$.
Unsurprisingly, we found that the model is per se well-calibrated at the token level. In this work, we will primarily focus on line-level completion. While several datasets exist for this problem (allamanis2013mining, 2, 51), we use DyPyBench (bajaj_piyush_2023_7886366, 6, 7), a new dataset, consisting of 50 popular open-source Python projects, including test suites for these projects. The test suites allow a test-correctness measure, in addition to the highly restrictive exact-match (viz., the original line). Figure 3b presents a sample prompt and model output for the line completion task, where we provide the model with all the prior lines and ask it to generate the next line.
DyPyBench consists of complex real-world projects, each with hundreds of thousands of lines of Python code, and totaling over 2.2 million lines of Python code [^7] As per naive sum of lines in all non-virtualenv *.py files across all project directories. We ran all test suites for each project with coverage reporting enabled. We subsequently extracted all functions from the projects using our derivative of function_parser introduced by (codesearchnet, 22), explored the dataset to ensure a lack of duplicates, and then joined each function with its line-level coverage report data. To make evaluation feasible, we randomly selected 1,988 functions with a body line span greater than 3, $100\%$ test coverage, and at least one line in the “Docstring”.
Docstring are defined in PEP257888PEP 257 – Docstring Conventions as a string literal that occurs as the first statement in a function/method definition. They frequently contain the function’s documentation and interface, and thus capture usage & intent. We then randomly chose a non-whitespace, non-comment line to be replaced (replacement line) for each function, which we reused across experiments. Finally, using a Python parser, we produced a variant of this dataset with all Docstrings removed, which we refer to as No Docstring. During our experiments, the model is provided the lines preceding the replacement line as the prompt, up to and including the signature. With this approach, we can examine how model calibration & completion quality is affected by natural language rich context i.e., documentation. Details of this pipeline are discussed in § 3.4.
3.2.3. Program Repair
Program repair is a well-studied problem in Software Engineering (goues2019automated, 18). Several studies have applied LLMs to the program repair problem and shown that it is effective (fan2023automated, 17, 26, 1). However, LLM calibration for program repair is not well-understood. Several program-repair datasets are available, including the widely-used Defects4J (just2014defects4j, 28), which includes real-world examples of buggy programs, with fixes and test-sets. This paper primarily focuses on line-level bug repair, represented-by 120 samples from the Defects4J dataset. However, with only 120 samples, we may not obtain a comprehensive view of calibration. Therefore, we included another dataset, ManySStubs4J (karampatsis2020often, 31) (abbr. SStubs), which consists of single-line repairs. Considering resource constraints, we selected 3,000 samples from this dataset, which is sufficient to investigate calibration. However, ManySStubs4J does not provide test-sets; so the only evaluation metric available is the exact-matching of the generated text to the ground truth bug-free text.
3.3. The models
3.3.1. Codex
We used three models for our experiments; OpenAI Code-davinci-002, which we refer to as Codex, is one of them. Codex is a family of GPT-3 based high-performance code models (chen2021evaluating, 12), which is used in GitHub Copilot products. The Codex is trained on large volumes of both natural language and source code from publicly available sources, including GitHub. Codex is most capable in Python, but it is proficient in over a dozen languages, including JavaScript, Go, Perl, PHP, Ruby, Swift, TypeScript, and even Shell [^9] https://openai.com/blog/openai-codex. We use Code-davinci-002 in our experiments, as it is the most capable version of the Codex family, with 175 billion parameters trained on a more recent text and code corpus compared to previous members of the family. It is a completion model; we access it using an API key. OpenAI provides logits for each of the generated tokens, from which we calculate token probabilities (See  § 3.1.2).
3.3.2. GPT-3.5-Turbo-Instruct
OpenAI GPT-3.5-Turbo-Instruct (abbr. GPT-3.5) is an instruction-tuned model trained to align with user intent across many tasks.
Unlike GPT-3.5-Turbo, GPT-3.5-Turbo-Instruct is not optimized for chat. Rather, it is optimized for directly answering questions or generating text. It offers similar compute performance to GPT-3.5-Turbo. As of the time of our experiments, chat models from OpenAI do not provide logits, and so GPT-3.5-Turbo and GPT-4 models could not be used for our calibration experiments. When available, we get token log probabilities using the OpenAI completions API. [^10] https://platform.openai.com/docs/api-reference/completions/create
3.3.3. CodeGen2
We also studied an open-source model called CodeGen2 (nijkamp2023codegen2, 45). CodeGen2 is an enhanced version of the CodeGen (nijkamp2022codegen, 46) model, and its authors have aimed to improve code synthesis performance by unifying four key components: (i) model architectures, (ii) learning methods, (iii) infill sampling, and (iv) data distributions. CodeGen2 is available in various sizes (1B, 3.7B, 7B, and 16B parameters), but we specifically utilized the largest and best performing 16-billion-parameter variant in our experiments. However, even the largest CodeGen2 model is much smaller than the OpenAI models we examine in this paper.
3.4. Summary of Complete Pipeline
Figure 4. Overview of our solving and evaluation pipeline
The pipeline consists of two core operations—code generation and code evaluation. Both are relatively independent, and both compute-intensive; the former requires either heavy GPU usage, or accessing the OpenAI API endpoints, and the latter requires checking generated code for syntax errors, and then running a large number of tests on a variety of different generations from different models. We leveraged this independence by querying the LLMs, and running the tests where available, in a parallelized environment. Each thread is responsible for utilizing one of the above LLMs to address a coding prompt and subsequently evaluate the generated code. The steps are given below, referring to Figure 4.
  1. (1)
    Solving: LLM Code Generation: The LLM is invoked to generate a solution for the coding prompt: Python code for DyPyBench, HumanEval, and MBPP datasets, and Java for the Defects4J and SStubs datasets. This step either uses the OpenAI API, or (in the case of CodeGen2), a local A100 80GB GPU. Post-solution, confidence measures are computed for the generated code which are used in our analysis. For patching & line completion tasks, all reference solutions are $<80$ tokens (median $\approx 10$ tokens). For function synthesis tasks, all references are $<440$ tokens (median of $\approx 40$ tokens, p90 of $\approx 115$ tokens). Therefore, we choose to sample up to a maximum of 500 tokens. We sample from the model with temperature=0, consistent with the reality that busy developers typically look at just the first suggestion in the completion task  (Barke_James_Polikarpova_2022, 8). For function synthesis task, temperature zero is most accurate and is fairly standard practice when doing pass@1 with only one solution to generate and present  (chen2021evaluating, 12, 39). In the case of function synthesis problems, we stop generation at the end of the prompted function block. In the case of line-level problems, we stop at a newline character. The completion is then merged into its original context. For example, in DyPyBench, the generated line replaces only the original line in the file, the project’s test suite is run in the DyPyBench Docker environment, ultimately producing a test report which is parsed and recorded.
  2. (2)
    Evaluating Correctness: Testing & Matching If generated code passes syntax checks, tests are run, and passing rate is evaluated; otherwise, all tests are marked as failing. For DyPyBench, SStubs, and Defects4J we also evaluated correctness using exact-match with the known version found in the dataset (exact match not shown in Table 3).
  3. (3)
    Analyzing: We record outcomes, including both the solution and its confidence measures; we then analyze and summarize them as presented in the results section.
4. Results
We begin with a brief overview of the findings on the correctness- & confidence- measures of LLMs on the various tasks, and then provide detailed results on the calibration-related research questions.
All Pass@1 Exact-Match
CodeGen2
Codex
GPT-3.5 CodeGen2
Codex
GPT-3.5
SStubs -
- 0.73% 27.77% 20.27%
DyPyBench 28.84%
32.96%
33.22% 19.68% 23.60% 23.96%
Defects4J 0.00% 23.33% 19.17% 0.00% 19.17% 15.00%
HumanEval 23.17%
47.24%
64.60% - - -
MBPP 29.08%
61.79%
72.04% - - -
Table 2. Performance comparison of models on tasks. Metrics are All Pass at Rank 1 (All Pass@1), meaning all project test cases passed with the line completion on first and only sample (at $t=0$), and Exact-Match, meaning the line completion was an exact string match with the original project line. Exact-Match is not commonly used for function synthesis tasks, since the generated output is longer and less likely to match. SStubs dataset does not have test cases. Boldface signifies high performing model for task and metric.
4.1. Descriptive Analysis of Results
All Pass@1 Exact-Match
Docstring
No Docstring
Docstring No Docstring
CodeGen2 28.84
23.10
19.68 15.10
Codex 32.96
21.69
23.60 17.77
GPT-3.5 33.22 26.77 23.96 18.72
Table 3. Performance comparison of models on DyPyBench Line Completion task: experiments include With and Without Docstring. Metrics are All Pass at Rank 1 (All Pass@1), meaning all project test cases passed with the line completion on first and only sample (at $t=0$), and Exact-Match, meaning the line completion was an exact string match with the original project line. Boldface signifies high performing model for dataset and metric.
4.1.1. Overall Correctness
Correctness performance rate of the various models on the various tasks and datasets, are presented in Table 2. Specifically, we report the fraction of samples passing all test cases for a given model and dataset, and the percentage of exact-matches. We found that GPT-3.5-Turbo-Instruct worked well for both function synthesis (HumanEval and MBPP), and line-level code completion (DyPyBench), whilst Codex generally performed well on program repair (SStubs and Defects4J). The DyPyBench benchmark reflects in some ways the most popular use of LLMs, viz., for code completion. In our experiments, we evaluated the line-level code completion test samples in DyPyBench both with and without the Docstring provided with the function (thus possibly relying on as little as the signature).
Table 3 compares the performance of the various models on the DyPyBench dataset, with and without the Docstring comment. Performance with the Docstring is substantially higher than without. The smaller CodeGen2 model shows the most gain in test-passing performance; the larger models show gains in both test-passing and exact-match, perhaps reflecting their greater capacity for recall with Docstrings, which lead to better in-distribution priors. However, as depicted in 4(a) and 4(b), it is evident that while the correctness measure improves with the presence of a Docstring, the version without a Docstring remains well-calibrated. This observation suggests that calibration cannot be solely inferred from the correctness measure.
4.1.2. Correctness: Test-passing vs. Exact-match
As per § 3.1.1, we evaluate correctness both on test-passing and exact-match. Our experiment included two datasets (Defects4J and DyPyBench), where both methods of measuring correctness were available. Since Defects4J consists of only 120 samples, we present the results for DyPyBench; in this case, as per Table 3, GPT-3.5-Turbo-Instruct performed best, with approximately 33% of generated code passing all available tests, and approximately 24% matching exactly. In this setting (DyPyBench/GPT-3.5) we cross-tabulate the performance across these two correctness-measuring methods in Table 4.
All Pass@1
Exact-Match False True Total
False 59.89% 16.16% 76.04%
True 6.89% 17.06% 23.96%
Total 66.78% 33.22% 100.00%
Table 4. Matrix of all tests passed versus exact-match notions of correctness for DyPyBench (with Docstring). Metrics are All Pass at Rank 1 (All Pass@1), meaning all project test cases passed with the line completion on first and only sample (at $t=0$), and Exact-Match, meaning the line completion was an exact string match with the original project line.
It is interesting to note that approximately half the test-passing generations did not match the original code exactly; furthermore 6.89% of the cases where the code matched exactly, did not pass all the test cases. Upon careful study we found that these tests were “flaky”, depending on network conditions, execution order, and other variable execution environment conditions. This aligns with (bajaj2023, 7), the author of this dataset, who observed an overall $8.07\%$ failure rate, but noted that 31 out of 50 projects had zero failed tests. This table illustrates the relative merits/demerits of each correctness-evaluating approach, in practical SE settings. Since the correctness performance is different with these two notions of correctness, the calibration is also different, as we see below.
4.1.3. Confidence Measures
As per § 3.1.2 we use two intrinsic measures of confidence: Average token probability $p_{\text{avg}}$, total generated sequence probability $p_{tot}$; and three reflective measures: verbalized self-ask $p_{v}$, a probability-derived confidence $p_{B}$ where we compute the probability on the model’s TRUE answer token, and finally a combined $p_{NB}$ variant which computes normalized confidence using the probabilities of the TRUE token and the FALSE token. We also provide a baseline confidence measure that uses the raw length $L$ of the generated sequence, linearly scaled such that solutions as short as the shortest reference solutions are 100% confident, and solutions as long as the longest reference solution are 0% confident, so that longer solutions are less confident. As might be expected, the two intrinsic measures $p_{avg}$ $p_{tot}$ are usually somewhat, and sometimes strongly, positively correlated with each other within the same model, dataset, and task; the two logit-based reflective measures $p_{B}$ $p_{NB}$ are usually strongly positively correlated, since one is calculated from the other. The reflective verbalized self-ask confidence measure $p_{v}$, and the two logit-based reflective confidence measures have no consistent relationships with other reflective measures, and vary by model, task, and dataset.
The intrinsic vs reflective measure values have no consistent relationship, even for a given model, dataset and task. This lack of relationship may not necessarily be negative: e.g., perhaps the model’s reflective, prompted confidence may be better calibrated, as suggested by prior work (bai2022constitutional, 5, 4).
4.2. Calibration Analysis
Line Completion Function Synthesis Program Repair
DyPyBench
HumanEval MBPP Defects4J SStubs
Model Metric
${\mathcal{B}}$
[…]
GPT-3.5 Total Prob 0.23 -0.03 0.15 0.62 -1.70 0.63 0.71 -2.50 0.71 0.25 -0.63 0.28 0.24 -0.50 0.25
Avg Prob
0.41
-0.87 0.46 0.27 -0.18 0.23 0.22 -0.09 0.14 0.68 -3.39 0.73 0.64 -2.94 0.69
Ask T/F
0.25
-0.13 0.16 0.34 -0.47 0.37 0.33 -0.64 0.38 0.15 +0.05 0.04 0.16 -0.01 0.04
Ask T/F N
0.25
-0.15 0.15 0.23 +0.01 0.19 0.22 -0.11 0.16 0.20 -0.30 0.24 0.22 -0.34 0.23
Verbalize
0.43
-0.92 0.42 0.28 -0.24 0.22 0.24 -0.17 0.17 0.58 -2.72 0.60 0.50 -2.09 0.53
Length
0.44
-0.99 0.46 0.23 -0.03 0.15 0.22 -0.10 0.16 0.53 -2.43 0.60 0.53 -2.26 0.60
Unskilled
0.22
0.00 0.00 0.23 0.00 0.00 0.20 0.00 0.00 0.15 0.00 0.00 0.16 0.00 0.00
[dashed] Codex Total Prob 0.23 -0.02 0.16 0.44 -0.77 0.45 0.60 -1.52 0.60 0.25 -0.39 0.24 0.20 0.00 0.09
Avg Prob
0.46
-1.07 0.50 0.34 -0.38 0.35 0.24 -0.03 0.19 0.66 -2.68 0.69 0.58 -1.90 0.62
Ask T/F
0.24
-0.09 0.12 0.37 -0.47 0.36 0.49 -1.06 0.50 0.18 +0.01 0.07 0.19 +0.03 0.02
Ask T/F N
0.23
-0.06 0.07 0.32 -0.29 0.30 0.42 -0.79 0.43 0.25 -0.41 0.27 0.23 -0.14 0.18
Verbalize
0.38
-0.74 0.35 0.42 -0.67 0.40 0.38 -0.61 0.33 0.47 -1.65 0.50 0.43 -1.14 0.42
Length
0.43
-0.95 0.45 0.44 -0.77 0.44 0.56 -1.35 0.56 0.50 -1.79 0.55 0.56 -1.78 0.59
Unskilled
0.22
0.00 0.00 0.25 0.00 0.00 0.24 0.00 0.00 0.18 0.00 0.00 0.20 0.00 0.00
[dashed] CodeGen2 Total Prob 0.21 -0.02 0.15 0.23 -0.30 0.23 0.29 -0.41 0.29 - - - - - -
Avg Prob
0.44
-1.16 0.50 0.60 -2.39 0.66 0.58 -1.80 0.61 - - - - - -
Ask T/F
0.23
-0.10 0.14 0.25 -0.38 0.24 0.25 -0.19 0.21 - - - - - -
Ask T/F N
0.33
-0.59 0.35 0.39 -1.19 0.45 0.39 -0.88 0.43 - - - - - -
Verbalize
0.42
-1.04 0.41 0.43 -1.39 0.42 0.40 -0.94 0.38 - - - - - -
Length
0.47
-1.28 0.51 0.38 -1.14 0.42 0.33 -0.58 0.28 - - - - - -
Unskilled
0.21
0.00 0.00 0.18 0.00 0.00 0.21 0.00 0.00 - - - - - -
Table 5. Calibration measured as raw, non-scaled Brier Score (${\mathcal{B}}$), Skill Score ($SS$), and Expected Calibration Error ($ECE$), with respect to “all passed” notion of correctness, except SStubs which is “exact-match”. CodeGen2 repair values are omitted as it does not perform the task with greater than 1% accuracy. The “Unskilled” row corresponds to a naive approach where the confidence is always returned as the base correctness rate, with Skill Score ($SS$) always zero by definition.
4.2.1. Calibration without Rescaling
Table 5 presents the results for Line Completion, Function Synthesis, and Program Repair for each model and the raw confidence measure, sans any rescaling method. We find raw confidence measures poorly calibrated, with inconsistent exceptions. In fact, the raw baseline rate (base rate of error) is hard to beat; the best skill-score is around $0.05$
Line Completion
Using the $p_{tot}$, confidence measure is slightly worse than the baseline rate; calibration error is modest ($ECE\sim 0.15$).
The total probability improves on the average probability, which is overconfident: the average token probability exceeds the $\sim 30\%$ overall success rate.
Function Synthesis
For function synthesis with raw measures, the Ask T/F normalized ($p_{NB}$) shows best calibration for GPT-3.5, slightly better than Unskilled, for HumanEval; the best for MBPP is $p_{avg}$. The intrinsic measures are inconsistent; with average probability showing indicators of good calibration for GPT-3.5 and Codex, but not for CodeGen2.
Program Repair
For program repair, we observe the strongest best-case performance with regards to the metrics. Both GPT-3.5 and Codex show positive $SS$ and low $ECE$ for Ask T/F ($p_{B}$), but they are inconsistent after normalization ($p_{NB}$). These metrics suggest with reflection, these models’ confidence is calibrated regarding repair correctness; however, the reliability plots (see appendix Figure 6), do not indicate good calibration on this task, from any confidence measure.
There are several caveats here. First, Defects4J is a small dataset, findings may not generalize. Second, CodeGen2 performs poorly on Defects4J. Since CodeGen2 is a smaller model without instruction tuning and relatively more limited reasoning capabilities it gets “distracted” by the buggy version shown in the prompt: it often just repeats the buggy lines. With very few correct outputs, the estimation of the confidence measure becomes unreliable. Therefore, we have removed the CodeGen2 results from Table 5. A final caveat is that SStubs uses only exact-match as a correctness measure, which is quite different from a test passing measure (see § 4.2.6). We now discuss results after applying rescaling.
4.2.2. Results with Rescaling
Table 6 shows the results after applying Platt scaling to all measures (See § 3.1.4). 4(c) shows a reliability plot before rescaling, and its equivalent after rescaling in 4(a). Rescaling can improve calibration. Considering all values, $ECE$ improves from an average of 0.32 to 0.03. For just those measures with a post-scaling $SS$ of at least 0.05, $ECE$ improves from 0.38 to 0.05.
(a) DyPyBench with Docstring, Platt scaled reliability plot
(b) DyPyBench without Docstring, Platt scaled reliability plot
(c) DyPyBench with Docstring, nonscaled reliability plot
(d) DyPyBench with
Docstring, ROC curve
plot
Figure 5. Reliability and ROC plots for DyPyBench line-level code completion tasks, with respect to All Pass @1 correctness measure and Average Token Probability confidence measure. GPT-3.5TurboInstruct was used for both experiments. Numbers at base of bars denotes number of samples in bin. $\mathcal{B}_{ref}$ refers to the unskilled predictor Brier, $ECE$ to Expected Calibration Error, $\mathcal{B}$ to Brier Score, and $SS$ to Skill Score. Red & purple lines represent scaled & non-scaled quantile bins rather than evenly spaced bins.
Understanding “Bucket Collapse”
Applying Platt scaling, however, can be deceptive. If a confidence measure is uninformative, as indicated by an AUC-ROC near 0.5, Platt-scaling results in all values being squished to the baseline rate. All values will thus be placed in a single bucket whose probability exactly matches the baseline rate, resulting in an $ECE$ near 0. This indicates a flaw of $ECE$. $SS$ and Brier, on the other hand, would reveal the poor calibration.
Thus when applying rescaling, it is important to look at Skill Score (or AUC), rather than just Brier and $ECE$.
Line Completion
After applying rescaling, only the intrinsic measures show an improvement in skill over the baseline rate. $p_{avg}$ and $p_{tot}$ are roughly equally calibrated. The calibration and skill appears roughly consistent between all three models in this case.
Function Synthesis
Rescaling improves calibration results for function synthesis. The $p_{tot}$ measure reaches a $SS$ of 0.18 for HumanEval. Additionally, similar $SS$ and $ECE$ values are found in GPT-3.5 reflective Ask T/F ($p_{B}$.
Program Repair
Rescaling doesn’t improve skill score, for any measure. As noted in § 4.2.1, there are limitations that make clear program repair conclusions difficult.
Line Completion Function Synthesis Program Repair
DyPyBench
HumanEval MBPP Defects4J SStubs
Model Metric
${\mathcal{B}}$
[…]
GPT-3.5 Total Prob
0.21
+0.07 0.03 0.20 +0.15 0.09 0.19 +0.07 0.05 0.16 -0.05 0.16 +0.03
Avg Prob 0.20 +0.08 0.04 0.23 -0.02 0.20 0.00 0.16 -0.03 0.16 +0.02
Ask T/F
0.22
0.00 0.20 +0.12 0.11 0.18 +0.09 0.06 0.15 +0.05 0.16 0.00
Ask T/F N
0.22
0.00 0.20 +0.14 0.07 0.18 +0.11 0.04 0.15 +0.04 0.16 +0.01
Verbalize
0.22
0.00 0.24 -0.05 0.20 +0.02 0.17 -0.09 0.16 0.00
Length
0.22
0.00 0.24 -0.03 0.20 +0.01 0.16 -0.06 0.16 0.00
Unskilled
0.22
0.00 0.23 0.00 0.20 0.00 0.16 0.00 0.16 0.00
[dashed] Codex Total Prob 0.20 +0.09 0.03 0.22 +0.11 0.08 0.22 +0.06 0.04 0.18 -0.01 0.19 +0.05 0.02
Avg Prob
0.20
+0.09 0.04 0.22 +0.14 0.07 0.21 +0.12 0.06 0.18 -0.02 0.19 +0.05 0.02
Ask T/F
0.22
0.00 0.24 +0.03 0.24 0.00 0.18 0.00 0.19 +0.04
Ask T/F N
0.22
0.00 0.24 +0.03 0.24 0.00 0.18 -0.01 0.20 +0.02
Verbalize
0.22
0.00 0.26 -0.02 0.24 -0.01 0.18 0.00 0.20 0.00
Length
0.22
+0.01 0.26 -0.03 0.24 -0.01 0.20 -0.10 0.20 0.00
Unskilled
0.22
0.00 0.25 0.00 0.24 0.00 0.18 0.00 0.20 0.00
[dashed] CodeGen2 Total Prob 0.19 +0.08 0.04 0.18 0.00 0.21 0.00 - - - -
Avg Prob
0.19
+0.07 0.02 0.17 +0.03 0.21 0.00 - - - -
Ask T/F
0.21
0.00 0.18 -0.01 0.20 +0.01 - - - -
Ask T/F N
0.21
0.00 0.17 +0.04 0.21 0.00 - - - -
Verbalize
0.21
0.00 0.18 -0.01 0.21 0.00 - - - -
Length
0.21
0.00 0.18 -0.02 0.21 0.00 - - - -
Unskilled
0.21
0.00 0.18 0.00 0.21 0.00 - - - -
Table 6. Calibration measured as Platt-scaled Brier Score (${\mathcal{B}}$), Skill Score ($SS$), and Expected Calibration Error ($ECE$), with respect to “all passed” notion of correctness, except SStubs which is “exact-match”. In cases where the $SS$ is less than 0.05, the $ECE$ is omitted. This is because an estimate without any signal will become Platt-scaled to approximately the base rate. This will appear as one well calibrated bin, resulting in an $ECE$ near zero, but does not provide information. CodeGen2 repair values are omitted as it does not perform the task with greater than 1% accuracy.
4.2.3. Is Rescaling a Panacea for Calibration?
As discussed in § 4.2.2, there is a clear trend that rescaling does improve calibration (RQ5). It has been used in settings other than generative models of code, with other notions of correctness  (guo2017calibration, 20, 43, 16, 49, 13, 37, 36). However, there are disadvantages. First, (see  § 4.2.2), “bucket collapse” can mislead with deceptively low ECE. Second, some correctness data is needed, to fit rescaling parameters.
These reasons suggest one must be careful when analyzing and reporting calibration results based on rescaling, and highlights the need for further work on confidence measures that might more directly calibrated.
4.2.4. RQ2: How well do intrinsic calibration approaches work?
Here we directly address RQ2 to see what conclusions we can draw from intrinsic measures. Without rescaling, total probability $p_{\text{total}}$ shows indications of calibration. With rescaling, there is a possible 10-20% improvement over baseline rates and good calibration, but it is inconsistent as skill is poor for CodeGen2 on Function Synthesis. As discussed later in § 4.2.6, results may vary significantly for exact-match correctness.
4.2.5. RQ3: How well do reflective calibration approaches work?
Without rescaling, results are inconsistent. In some cases, such as GPT-3.5 HumanEval and Defects4J, there are signs of calibration with $SS$ values of reaching 0.01 and 0.05 respectively. Normalizing the T/F values induces a notable difference; but there are inconsistencies vis-a-vais tasks and models. For nonscaled GPT-3.5 results, $p_{NB}$ improves calibration in Line Completion and Function Synthesis by an average of -0.34 $SS$ and 0.19 ECE, but not for Program Repair or for CodeGen2. Rescaling generally removes any sign of a normalization trend. For the alternative reflective approach of verbalization, the probability is not well calibrated for these models on the studied SE tasks. In some cases, the reflective approaches are best calibrated without scaling (see Table 5).
Alternative and improved ways of prompting (e.g., few-shotting (Kadavath2022LanguageM, 29), different verbalization formats (lin2022teaching, 38, 58), fine-tuning (Kadavath2022LanguageM, 29, 38), chain-of-thought (Wei2022ChainOT, 61, 58), etc.) may alter these findings and are areas for future work.
4.2.6. RQ4: How calibrated are LLMs for code with respect to “correctness-modulo-testing” and “exact-match” notions of correctness?
We discuss this question with respect to DyPyBench, since it is a realistic, large-sample dataset supporting both notions of correctness. As discussed earlier (§ 4.1.2), OpenAI models perform better overall (regarding correctness) on test-passing than on exact-match; but they are better calibrated for exact-match! Here, the OpenAI models’ intrinsic $p_{tot}$ probability exhibits very high skill for predicting exact-match correctness, in the range of 0.31 to 0.43 (appendix Table 7). The same Codex $p_{tot}$ shows the highest skill for test-passing correctness, but with much lower Skill Scores (around 0.09). Thus the measure can skillfully predict when it repeats the exact sequence of human-written code found in the corpus, but it is less skilled in predicting when the code is correctly test-passing. This kind of “data recovery” could be expected, given the original training objective of LLMs (Carlini2020ExtractingTD, 11). It is also noteworthy that the exact-match skill for the same models for the SStubs dataset is lower (max ~ 0.06), possibly because these are cases where the human programmer originally made mistakes, and as reported by (10174227, 25, ), Codex tends to prefer to repeat those errors, rather than their fixes. Thus the exact-match correctness in SStubs is more demanding, because the desired code in this bug-fix dataset is different from the originally human-written versions.
In conclusion, a measure well-calibrated for exact-match will inform the user when the model is likely producing code similar to the human-written code has seen in training; this may not always be desirable, since the training data may include bugs. If a confidence measure can be found that is well-calibrated for test-passing, especially if such a measure is trained on code with good test-sets, it may be preferable.
5. Discussion
Language models are now widely-integrated into Software Engineering practice, via tools like Copilot (ziegler2022productivity, 71) and Didact (researchLargeSequence, 35). We raise here the importance of calibration when integrating generative LLMs in coding practice. We evaluate the calibration of generative LLM use (especially code completion) with large samples of realistic data (DyPyBench, SStubs), using widely adopted models, as well as some more academic datasets (HumanEval, MBPP).
Using a well-calibrated model–beyond simple defect prediction
To clarify how a well-calibrated model enables more well-grounded decision-making concerning generated outputs, as compared to a traditional ROC curve—we consider Codex working on code completion, where correctness is determined by testing passing, and confidence is measured by $p_{avg}$, average token probability. The base correctness (test-passing) rate of completions is about 33%. We find here that re-scaled $p_{avg}$ has a Skill Score of 0.09 and $ECE$ of 0.04.
If using the ROC curve instead, (see 4(d)), one might cautiously choose a high-confidence threshold such that the FP rate is very low (of course TP rate would be low as well) and thus avoid having to review much code (while also regrettably not receiving as much help from the LLM). However, this measure $p_{avg}$ is well-calibrated (see  4(a)) allowing a more graduated set of decisions. This diagram is well calibrated up to around 90% confidence, above which it becomes over-confident; fortunately, there are relatively few samples (around 1.5% of the data) at this high level. Therefore, for most of the confidence scale, a user could simply look at the confidence level, and get a very good idea of how likely the code is to be correct, and make a graduated set of decisions about the increasing level of scrutiny required as the model’s confidence decreases. This provides an illustration of the greater benefit provided a well-calibrated measure (high Skill level, low ECE) over one that is just providing a good AUC; the latter does not allow such graduated deployment of quality-control effort.
User studies, and perhaps training, would be helpful, as developers learn to use calibrated probabilities in decision-making.
Beyond simple “correctness”
In addition to the above uses, which considered a single notion of “correctness”, one could consider a multi-class correctness prediction task, where the model could indicate the confidence in correctness (the absence of defects) from multiple perspectives: severity of possible defect, the kind of defect (relating to security, integrity, privacy, fairness, etc) and defect complexity (indicating the cost or schedule impact of repairs). Drawing an analogy to classical forecasting, this is analogous to not just the probability it will rain, but probability it will be a drizzle or be a drenching thunderstorm.
Why calibration now?
We’ve always had bugs; poor-quality code isn’t new. Our push for calibration, however, arises from the increasing amount of code generated by LLM. GitHub claims that up to 61% of code [^11] https://github.blog/2023-02-14-github-copilot-now-has-a-better-ai-model-and-new-capabilities/ in some systems is generated by LLMs. It is also known that LLMs make a lot of mistakes. A recent paper has reported (10174227, 25) that LLMs, even when trained on properly fixed code, tends to recapitulate the the old unfixed, buggy code when prompted in context. However, LLMs do have very high capacity, and a demonstrated ability to usefully reflect (shinn2023reflexion, 54, 5) on their generated text. Thus we have both a high risk (of buggy code), and a chance to improve productivity. We believe that improved calibration could lead to better management of the risk-benefit of LLM-generated code. The studied correctness calibration is a stepping stone for more complex notions of confidence (like severity and localized confidence). Additionally, by studying code LLMs, we might make progress on the general safe deployment of capable generative models (gros2023ai, 19).
6. Threats to Validity
We now discuss threats to validity of our results.
Sample Size
While three of our datasets contain more than 900+ samples each, HumanEval and Defects4J datasets consist of only 164 and 120 samples, respectively. Results on these datasets may not generalize. However, we note that our study has a large and natural dataset for the line-level code completion task, which has current practical importance. Given the noise and variance we observe, we recommend future work push towards larger and more natural in particular for Function Synthesis & Program Repair.
Artificial vs. real world data
For function synthesis, we used the popular HumanEval and MBPP function synthesis datasets. However, these datasets contain fairly small Python programs that may not represent real-world software development functions. However, our other datasets, such as DyPyBench and SStubs are quite representative of real-world, open-source GitHub projects. As seen in section 4, the models were best calibrated (highest SS) on the function synthesis task; but results on this task/dataset may not generalize to real-world programming. However, we did see some consistency in the results, between the code completion task and the function synthesis task: the $SS$ was positive for intrinsic confidence measures with OpenAI models. This suggests the possibility for calibrated decision-making in the quality control of LLM-generated code.
Implementation Issues
While efforts were taken for careful implementation, possible issues in implementation is always a threat. A known issue includes that a small number (¡3/dataset-model) of CodeGen generations failed due to CUDA/OOM errors. A small number (¡4/dataset-model) of values also encountered unexpected exceptions when evaluating test cases.
Experimental Design
Our exploration is not exhaustive; other SE tasks and datasets also could benefit from calibration studies. Additionally, the specific prompts we used for this paper surely played a role in our findings. Other prompts or problem phrasings (such as different forms of context for line-level code completion) may yield different results. Regarding test flakiness: the test “flake” rate of DyPyBench is not zero, but is not unrealistic (luo2014empirical, 42). Additionally, conclusions may shift as models continue to scale and improve, with possibly emergent capabilities.
Still, our study, which includes three tasks and five datasets, provides a good starting point for further studies.
7. Related Work
LLMs for code are extensively studied (Zhang2023ASO, 67, 69). While calibration has a long history in modeling (brier1950verification, 9, 56), it is not a frequently studied topic in the SE community. Early work moving into modern machine learning studied the calibration of smaller neural models performing classification tasks on text and images; while these early models were poorly calibrated per se, their performance could be improved by simple scaling (guo2017calibration, 20) of their output probabilities. As models became larger, calibration was found to improve (srivastava2022beyond, 55). Pre-training was also found to improve calibration (hendrycks2019using, 21, 16); however, these findings have been disputed (chen-etal-2023-close, 14).
More recent works evaluated LLM calibration on a wide variety of settings (kadavath2022language, 30, 27, 16, 32).  (desai2020calibration, 16, ) studied non-code (natural language) tasks such as inference or paraphrasing, with only intrinsic measures using older-generation models (BERT and RoBERTA). (jiang2021can, 27, ) studied calibration for natural language question-answering using just intrinsic measures. In contrast, we study calibration for three coding-related tasks, using both artificial and natural code datasets, and both intrinsic and reflective confidence measures, to evaluate calibration in the SE domain.
(Li2019OperationalCD, 36, ) investigate the calibration of Computer vision (CV) models from an operational perspective i.e., the shift between training input and production inputs, presenting it as a software quality problem that can be addressed using Bayesian approaches. (Minderer2021RevisitingTC, 43, ) evaluate the calibration of at the time, state of the art CV models and find improved calibration with more recent models, notably those not using convolutions. (Park2022OnTC, 49, ) study the effect of the mixup technique (Zhang_Cisse_Dauphin_Lopez, 66) on calibration in a natural language understanding (NLU) setting using older generation models (BERT and RoBERTa). (Chen_Yuan_Cui_Liu_Ji_2023, 13, ) investigate the calibration of pretrained language models on various NLP tasks, also using older generation models (RoBERTa and T5). (Liang2023HolisticEO, 37, ) introduce the HELM benchmark, which includes calibration as one of its seven metrics to evaluate language models in a natural language context.
When suitably prompted, (kadavath2022language, 30, ) found that LLMs can output well-calibrated scores on whether their own answers are correct or not, viz., larger models “know what they know”. While this work did investigate some function synthesis tasks (HumanEval & an unpublished Python function dataset), they did so using only their private models, and ultimately focused on natural language tasks. (Key_Li_Ellis_2022, 32, ) developed an approach that given a natural language problem description, produces a confidence score for a sampled candidate solution based on generated specifications, allowing them to judge whether the LLM can solve the problem at all. Their metrics include calibration. Recent work has also explored calibration of software topics such as root cause analysis (zhang2023pace, 65).
8. Conclusion
In this paper, we argue that calibration would help better integration of LLM-generated code into software processes. We compared various measures of calibration, and argue that one should examine a combination of Skill Score (or AUC if scaling) paired with $ECE$ when observing a positive SS. We studied the calibration of intrinsic and reflective confidence measures in several practical settings (completion and repair) and a widely-used competitive setting (synthesis), and discuss the potential for better process integration of LLMs. We find that large models are generally poorly calibrated on code synthesis tasks, across a wide variety of confidence measures, with an average Skill Score of -1.60. We also find that Ask T/F has the highest average Skill Score of -1.60, suggesting that without access to rescaling, asking the model to self-evaluate and examining the response logit(s) may be the best option. Despite poor calibration a priori, we find that in some cases [^12] Generally, Total & Average Token Probability, Ask T/F and its normalized form exhibit positive skill i.e. above base rate after Platt scaling, Platt scaling results in well calibrated confidence measures, for example, GPT-3.5 Total Token Probability has an average raw Skill Score of -1.07 and $ECE$ 0.41, but the single highest scaled Skill Score across all experiments of 0.15, resulting in scaled Total Token Probability also having the highest average Skill Score of 0.05 (which is coincidentally the same number for average ECE).
Taking these findings into account, our work suggests the SE community should apply a rescaling approach, such as Platt scaling discussed in this work, on the following confidence measures: Total & Average Token Probability, Ask T/F or its normalized form, to gain calibrated confidence measures for graduated decision making in AI4SE processes.
That being said, all explored confidence measures have issues. It is not clear any studied measure is “deployment ready”, and one must be mindful of robustness to varying scenarios. Our findings also suggest that calibration based on exact-matching may be high, but carries the risk of recapitulating existing errors; calibration based on test-passing is currently not as good, and should perhaps be further studied and improved. We hope that our work clarifies the role of calibration when using LLMs in the software process, and leads to better calibration techniques across LLM domains, and more flexible/adaptive process integration.
9. Acknowledgments
We acknowledge the Intelligence Advanced Research Projects Agency (IARPA) under contract W911NF20C0038 for partial support of this work, as well as the National Science Foundation, under CISE SHF MEDIUM 2107592. Our conclusions do not necessarily reflect the position or the policy of our sponsors and no official endorsement should be inferred.
Appendix
Line Completion Program Repair
DyPyBench
Defects4J SStubs
Model Metric
${\mathcal{B}}$
[…]
GPT-3.5 Total Prob 0.11 +0.38 0.03 0.13 +0.01 0.16 +0.03
Avg Prob
0.13
+0.31 0.03 0.13 +0.01 0.16 +0.02
Ask T/F
0.18
0.00 0.12 +0.05 0.16 0.00
Ask T/F N
0.18
0.00 0.13 +0.02 0.16 +0.01
Verbalize
0.18
0.00 0.13 -0.03 0.16 0.00
Length
0.17
+0.06 0.03 0.14 -0.06 0.16 0.00
Unskilled
0.18
0.00 0.13 0.00 0.16 0.00
[dashed] Codex Total Prob 0.10 +0.43 0.02 0.15 +0.01 0.19 +0.05 0.02
Avg Prob
0.12
+0.34 0.02 0.16 -0.01 0.19 +0.05 0.02
Ask T/F
0.18
+0.01 0.16 -0.01 0.19 +0.04
Ask T/F N
0.18
+0.02 0.16 -0.01 0.20 +0.02
Verbalize
0.18
0.00 0.16 -0.01 0.20 0.00
Length
0.16
+0.09 0.02 0.17 -0.11 0.20 0.00
Unskilled
0.18
0.00 0.16 0.00 0.20 0.00
[dashed] CodeGen2 Total Prob 0.09 +0.41 0.01 - - - -
Avg Prob
0.11
+0.30 0.01 - - - -
Ask T/F
0.16
0.00 - - - -
Ask T/F N
0.16
0.00 - - - -
Verbalize
0.16
0.00 - - - -
Length
0.15
+0.06 0.03 - - - -
Unskilled
0.16
0.00 - - - -
Table 7. Calibration measured as Platt-scaled Brier Score (${\mathcal{B}}$), Skill Score ($SS$), and Expected Calibration Error ($ECE$), with respect to “exact-match” (EM) notion of correctness, excluding function synthesis tasks as EM is not a useful or commonly used notion of correctness. In cases where the $SS$ is less than 0.05, the $ECE$ is omitted. This is because an estimate without any signal will become Platt-scaled to approximately the base rate. This will appear as one well calibrated bin, resulting in an $ECE$ near zero, but does not provide information. CodeGen2 repair values are omitted as it does not perform the task with greater than 1% accuracy.
Line Completion Function Synthesis Program Repair
Model Metric
DyPyBench
HumanEval MBPP Defects4J SStubs
GPT-3.5 Total Prob
0.67
0.77 0.70 0.54 0.61
Avg Prob
0.68
0.61 0.56 0.57 0.60
Ask T/F
0.54
0.73 0.71 0.67 0.54
Ask T/F N
0.53
0.74 0.73 0.67 0.57
Verbalize
0.53
0.54 0.60 0.43 0.51
Length
0.53
0.64 0.61 0.48 0.53
Unskilled
0.50
0.50 0.50 0.50 0.50
[dashed] Codex Total Prob
0.68
0.70 0.66 0.52 0.66
Avg Prob
0.68
0.71 0.69 0.57 0.65
Ask T/F
0.49
0.65 0.56 0.62 0.61
Ask T/F N
0.49
0.61 0.53 0.54 0.59
Verbalize
0.52
0.46 0.50 0.45 0.49
Length
0.55
0.51 0.50 0.43 0.49
Unskilled
0.50
0.50 0.50 0.50 0.50
[dashed] CodeGen2 Total Prob
0.68
0.44 0.52 - -
Avg Prob
0.66
0.67 0.52 - -
Ask T/F
0.52
0.39 0.59 - -
Ask T/F N
0.54
0.34 0.57 - -
Verbalize
0.51
0.49 0.51 - -
Length
0.55
0.39 0.47 - -
Unskilled
0.50
0.50 0.50 - -
Table 8. AUC-ROC score of each technique
Figure 6. Calibration plots per model, confidence measure, and task. The blue bars represent nonscaled, evenly spaced bins. The orange bars are Platt scaled bins. The red lines represent five points of equal count quantiles (an equivalent number of problems in each bin).
Figure 7. Reliability plots for GPT-3.5, from left to right: HumanEval Ask T/F (Scaled), MBPP Ask T/F (Scaled), and HumanEval Ask T/F (Nonscaled). Red line denotes five quantiles. All three examples have similar AUC (0.73, 0.71, 0.73) but vastly different $ECE$ (0.11, 0.06, 0.37).
Figure 8. AUC vs Scaled $ECE$ (left) and AUC vs Scaled Skill Score (right) for GPT-3.5 confidence measures on all tasks. Shows a limited relationship between AUC and ECE, but a strong relationship between AUC and scaled SS.
Understanding effects of verbalized retry failures
Our implementation for verbalized confidence prompts the model to output the probability its generation is correct at temperature 1.0. If it does not contain a probability, then we resample up to 3 times. If that loop fails, then a confidence of 0.5 is returned.
This implementation seemed reasonable at the time (if the model won’t tell you its confidence, just go with the maximum uncertainty of 50-50), but after collecting the data and analyzing results we reconsidered, as 50% values might be overrepresented in the data.
To try to estimate how this might have influenced our results and conclusions, we searched for instances where the verbalized confidence was 50% (this provides a upper bound on how often this happens. There can also be cases where the model actually verbalizes a 50% confidence). This is relatively rare for GPT-3.5 (with the mean dataset having 0.04 of instances as 50%, range 0.02-0.08). It is more common for Codex (mean 0.14, range 0.09-0.17) and for CodeGen2 (mean 0.10, range 0.09-0.12). This evidence of how the instruction tuned models are more likely to actually perform the prompted task.
We reran our analysis excluding all these instances. We do not believe a different handling of these fail-retry values would have greatly changed our conclusions. In the scaled case, the Skill Score on average did not change (mean diff of 0.00) with extreme change of -0.04 $SS$ when already low skill. In the nonscaled case there were some drops in calibration (mean $SS$ change of -0.11 and mean $ECE$ change of 0.02). The more extreme changes areas of already poor calibration.
It is not clear what is the best default is in the situation where the model fails to verbalize a probability. It is not particularly valid to exclude these instance. More exploration is needed on this and the effects.